OpenStudy (anonymous):

A system of equations is given below: y = 4x − 3 2x + 7y = 41 Which of the following steps could be used to solve by substitution? 2(y = 4x − 3) + 7y = 41 4x − 3 = 2x + 7y 2x + 7(4x − 3) = 41 2(y = 4x − 3)

OpenStudy (anonymous):

I will face and medal

OpenStudy (anonymous):

I don't understand. Can you type the separate steps options?

OpenStudy (anonymous):

2(y=4x-3)+7y=41 4x − 3 = 2x + 7y 2x + 7(4x − 3) = 41 4x − 3 = 2x + 7y 2x + 7(4x − 3) = 41

OpenStudy (anonymous):

2(y = 4x − 3) + 7y = 41 4x − 3 = 2x + 7y 2x + 7(4x − 3) = 41 2(y = 4x − 3)